# A Typology of Decision-Making Tasks for Visualization

### v1

@@ 4 Typology of Decision-Making Tasks
As we outlined in the previous section, we created a typology that reflects and captures the ubiquity of decision-making while allowing for flexibility in the execution of decisions. Here we define the tasks of our typology, focusing on the unique properties of each. We then move to a discussion of the composability of the tasks, followed by four case studies that illustrate the outlined properties in real-world visualization systems.
4.1 Decision-Making Tasks
Our typology consists of three tasks derived from the scientific literature[28, 27]: CHOOSE, ACTIVATE, and CREATE. Each task is a function that represents a specific and distinct decision problem. The type of the inputs to these functions does not change the core process of the decision task. Some of the key differences between the tasks are the unique transformations of inputs into outputs, the existence of input features for evaluation, and the existence of well-defined evaluation functions.
4.1.1 CHOOSE
The CHOOSE task assesses options and outputs a subset deemed optimal or best. Specifically, CHOOSE takes as input a set of objects, and an evaluation function that determines a score for each input. Importantly, the evaluation of one input depends on the other inputs. In other words, whether an input is kept or selected for output depends on how well it scores relative to the other inputs.
Consider the example decision of buying a car; we have four cars from which we can CHOOSE. We could score the cars on their features relative to the other cars in the set, and select the car with the top score. Subsequent updates to the scoring function and the re-selection of the best car are represented as multiple executions of the CHOOSE task.
The core component of CHOOSE—the evaluation of inputs with respect to others—does not change, but there is flexibility in how evaluation is performed. As we discussed above, one way to execute the evaluation in a CHOOSE decision is to score each input, sort the alternatives according to that score, and select the top $k$. Alternatively, the evaluation could perform comparisons between each pair of inputs, keeping only the subset that wins in each comparison, and continue this process until $k$ inputs are left in the subset. In this case, no scoring function is required as long as there is a way to compare the inputs.
4.1.2 ACTIVATE
The ACTIVATE task represents a decision where inputs are evaluated, and only those that meet or exceed a threshold are returned. As in the previous task, ACTIVATE determines a score for each input using existing attributes on the inputs. However, the evaluation of one input is independent of the other inputs. Whether an input is returned in the output depends only on how well it scores, regardless of the scores of the other inputs.
In the car-buying example, the ACTIVATE task can be represented by applying thresholds to each existing feature and buying the car that satisfies all thresholds. For example, we could buy the car that is under a particular price and mileage and has a cargo capacity over a particular value. Notice that because each car is evaluated independently, we can assess them all in the evaluation simultaneously. Subsequent updates to the thresholds to eliminate all but a single car are represented as multiple executions of the ACTIVATE task.
Similar to the previous task, there is flexibility in the evaluation that occurs in ACTIVATE. For example, the threshold can change across inputs as a result of the features of each individual input. In the car buying decision, the price threshold may decrease if the mileage and cargo capacity are too close to the thresholds, or the price threshold could increase if the cargo capacity significantly exceeds the threshold. Importantly, this change in threshold does not depend on the other inputs or their features. Unlike CHOOSE, where evaluation does not require a scoring function because comparison or preference could be used instead, the ACTIVATE task must have one. This is because in ACTIVATE there is no other way to assess whether an input is included in the output.
4.1.3 CREATE
The CREATE task represents decisions on assembling, synthesizing, or generating new information. Specifically, CREATE can take any number of inputs (including none) and return outputs of a new type or augment inputs with new information. This is in contrast with the other tasks in our typology, where the inputs and outputs are consistent with each other. The evaluation in CREATE addresses how to assemble the inputs into outputs, or what outputs to produce given the inputs.
Consider again the decision to buy a car. Rather than deciding on an existing set of cars, the CREATE decision allows for three possibilities: (1) the creation of a new option that is not in the current set of options, (2) the addition (or removal) of a feature, and (3) synthesizing new information based on the inputs.
Of the three possibilities, the third use of CREATE is a departure from most decision-making literature and models, but is of particular importance to the visualization community. This decision represents the creation of knowledge or insights[9, 18, 38, 55], or more tangibly, reports[19], new data transformations[37], etc.
Similar to the other two tasks, the evaluation in CREATE can change according to the assessment of the input features. For example, we could consider building multiple possible cars, where each new car we add to the set is as different as possible from the others. In this case, the output of the CREATE task, the set of cars, is also an input alongside the set of features from which to build a new car to add to the set.
5 Composability
The decision tasks in our typology can be composed of or decomposed into other tasks in the typology. This is an important property that allows us to represent the hierarchies and complex structures that exist in real-world decision-making problems. For example, consider the decision to buy a car as a CHOOSE task. We could decompose this task into several sub-decision tasks like a CREATE decision which adds new possible cars to choose from, an ACTIVATE that filters cars that do not meet specific requirements, and a CHOOSE decision to select the best option. The ability to break down the tasks in our typology allows us to represent how smaller decisions help resolve the overall decision. This example illustrates two properties of our typology:
Composability: The three sub-decision tasks are interconnected to form a larger decision task. These sub-decisions are said to be composable because the output of one sub-decision task becomes the input of the next. For example, CREATE adds a decision critera, which is then used by ACTIVATE to filter out undesirable options. The property of composability is important for describing visualization tasks in general, as demonstrated by the task typology by Brehmer et al.[13]. The composition of these decision tasks can be visualized as a directed graph, referred to as a “diagram” throughout this paper. In these diagrams, directed edges represent the compositions of decision tasks. Importantly, when two tasks are connected by a directed edge, the input type of the task at the end of the edge must match the output type of the task at the beginning of the edge.
Hierarchy: The decomposition of the CHOOSE task in the above example into three sub-decision tasks implicitly defines a hierarchical structure of decision tasks. These hierarchies can be useful as they can describe a decision task at a high-level abstraction, while also offering detailed step-by-step decision-making processes at lower levels. While these hierarchies can be constructed either from the top-down or the bottom-up, during our interview study, we observed a preference among participants for a top-down approach (see Section 7). The participants started with high-level decision tasks and recursively decomposed them into sub-tasks until they achieved the desired level of granularity. In the case studies below, we demonstrate how these decision-task hierarchies can be created through a series of decompositions.
@@ 6 Case Studies
To illustrate the composability and hierarchical properties of our typology, we present four case studies where our typology is used to describe user decisions in the context of decision-support tools.
@@ 6.2 Case Study 2: Umbra
Level 1: At the root level, Umbra supports the ACTIVATE decision where a sanitized dataset is being evaluated on whether the changes made to it are acceptable or not, depending on its remaining utility and the privacy achieved.
@@ 6.3 Case Study 3: Centaurus
Level 2: The Level 1 ACTIVATE decision is supported by additional decisions – an ACTIVATE to select relevant article based on search terms, a CHOOSE to choose articles for deeper analysis, and another ACTIVATE to search for similar articles after a user has made a decision about a potential article to investigate further.
@@ Discussion:
The WireVis case study diverges from previous examples in that the root decision is not to simply choose or filter out potential options. Instead, it involves the creation of insights assembled from evidence gathered in the tool. The CREATE decision deviates from the typical multi-criteria decision-making (MCDM) style, and is a common and important aspect of decision-making within the visual analytics community.
@@ 7 Evaluation
Figure 6: Diagrams created by our participants during the interview study. Texts, labels, and annotations that contain identifiable information have been removed to preserve anonymity.
We recruited nine visualization experts to evaluate our typology from their perspectives and identify its strengths and weaknesses. Three of our participants are assistant professors, four are scientists at national laboratories, one is a senior visualization developer at a market research company, and one is a PhD student in visualization.
The interviews were semi-structured with a set of pre-determined open-ended questions with the opportunity to explore particular discussions further[36]. The interviews were conducted remotely via Zoom, each lasting approximately one hour, and recorded with the consent of all the participants. We asked the participants to choose a decision support tool they developed or an academic paper they wrote about during the interview. Then, we demonstrated our typology and its ability to compose decision-making tasks. This was followed by an exercise in which they had to describe the problem their tool addressed using our typology. After the exercise, we asked the participants open-ended questions about the properties of the typology. We concluded by asking them to reflect on other typologies and taxonomies in visualization and any suggestions for future work.
Overall, our interviews examined three broad areas: the typology’s correctness, utility, and comparison with existing task taxonomies and typologies commonly used in the visualization community. We address each of these areas in the following subsections.
@@ 7.1.1 Completeness
In evaluating the correctness of our typology of decision-making tasks for visualization, we first assessed its completeness to determine if there were any missing decision tasks in our typology that could limit its ability to capture all aspects of decision-making. The participants in our study found that our typology was indeed comprehensive, allowing them to represent all decisions within the problem domain. As noted by one of the participants: “I think [the typology] is a pretty good set of high level operations. […] It took me a little bit to really clarify the distinction between activate and choose. But once I got there, it makes a lot of sense. I don’t know if I can really think of another type of operation that I felt was missing. It seemed, at least in the case of my problem, a good set [of operations] to describe what’s going on.” Other participants expressed similar sentiments. Overall, all of the participants were able to describe their decision-making problems using our typology, and none of them identified any missing decision tasks or fundamental operations in our typology.
Interestingly, a few participants incorporated additional annotations when describing their decision-support tools using our typology (e.g., see Figure 7). These annotations were used to detail the contexts of these participants’ decision-making problems, including temporal information (e.g., time duration for how long a task would take or how frequently a task needs to be performed), roles (e.g., who in an organizational hierarchy should make the decision. Or, in a collaborative setting, which collaborator is responsible for which task), and decision styles (e.g., rational, conceptual, analytical, etc.). Such contextual annotations complement our typology and suggest an interesting future extension of our work. We will explore these concepts in greater detail in Section 8.
@@ 7.1.2 Expressiveness
Participants unanimously expressed their proficiency in composing and decomposing decisions with the typology, enabling them to construct cohesive representations of the problems their respective visualization tools addressed. They noted that the typology adeptly captured the decision-making processes inherent in problem domains and facilitated the visualization of user interactions with decision-support tools. As one of the participants said, “[The use of the typology] really kind of helped highlight some of the complexity around the design goals that went into it. And it’s definitely reflective of conversations that [my team and I] had going into the design of the tool and sort of throughout its life span. So, in that sense, I think it’s very representative.

### v3

@@ 4 Typology of Decision-Making Tasks
Our typology consists of three tasks, which we derived from examining the scientific literature and iteratively refining the definitions to meet our design goals (see Section 3). These tasks reflect and capture the ubiquity of decision-making, flexibly describe the execution of decisions, and the definitions use language consistent with the decision-making literature. The tasks are agnostic to who or what performs the evaluation, which ensures the typology can describe human or computational decision-making problems. Furthermore, our task definitions are agnostic to the rationality of the evaluations, which allows for our typology to be consistent and used in conjunction with computational and cognitive biases[24, 45]. We illustrate each decision task visually in Figure A Typology of Decision-Making Tasks for Visualization.
Each task in our typology evaluates and returns options according to its definition. Options are information entities typically relevant for a decision task. Options can have features that describe characteristics of the options. Lastly, criteria are preferences or standards on features of the options. As an example, we could consider options as cars, with features of color, price, and mileage, and an example criteria is “any color except white".
4.1 [Uncaptioned image]
The CHOOSE task involves assessing a set of options and returning a subset deemed optimal or best. Fischhoff’s concept of decisions with fixed options, and Simon’s final decision-making stage, choice, form the basis of the CHOOSE task. We more specifically define CHOOSE as deciding on the top or best $k$ options available (illustrated in the left panel of Figure A Typology of Decision-Making Tasks for Visualization). Importantly, within CHOOSE, the evaluation of options is dependent on the other options. Whether an option is returned depends on how it compares with the other available options.
Consider the example decision of buying a car; we have four cars from which we can CHOOSE. Rather than looking for a car that satisfies a set of criteria, we are looking for the best car from the current set available. Furthermore, the notion of best can change over time, which can lead to different choices. These changes could be due to different environmental or circumstantial contexts (e.g., living in a new location), or shifting perceptions (e.g., brand loyalty). Additionally, cognitive biases like loss aversion or the framing effect [24], and machine biases like the linking bias [45], are examples of biased evaluations that can occur in the CHOOSE task.
Unlike the other tasks in our typology, with CHOOSE we know how many options are returned upon execution, meaning only one execution of the CHOOSE task is required to reach a subset of options of size $k$. Although criteria can inform comparison of options in CHOOSE, criteria are not necessary to execute CHOOSE. Furthermore, it is not a given that the returned options meet a set of criteria. In other words, we have no guarantees on the quality of the returned options, only that they are the best subset of all evaluated options.
4.2 [Uncaptioned image]
The ACTIVATE task draws inspiration from Fischhoff’s decisions about action thresholds but expands beyond merely temporal thresholds. This task represents a decision where options are evaluated, and only those that meet or exceed a threshold are returned (illustrated in the middle panel of Figure A Typology of Decision-Making Tasks for Visualization). In contrast to CHOOSE, the evaluation of one option is independent of the others. Whether an option is returned depends only on its evaluation individually, regardless of the evaluation of the other options.
In the car-buying example, the ACTIVATE task can be represented by applying thresholds to each car such that only cars under a particular price and mileage are left after evaluation. Notice that because each car is evaluated independently, we can assess them all in the evaluation simultaneously. There is flexibility in the evaluation that occurs in ACTIVATE, as thresholds need not be strictly enforced across options or well-defined. For example, in the car buying decision, a person may exclude a car that is very near (but still under) both the price and mileage limits, but keep a car that is slightly over the price limit because the mileage is much lower than the limit. Some example biases that can occur in this evaluation include the cognitive biases of information bias and the anchoring effect[24], or the algorithmic processing bias in computational decision-making[20].
Unlike CHOOSE, where we know the exact number of options that are returned, we do not have this guarantee for ACTIVATE. It is possible that none or all options are returned by ACTIVATE. We have to execute ACTIVATE an unknown number of times to reach a subset of options of size $k$. However, any options returned by ACTIVATE meet a level of acceptability that is informed by given criteria. In other words, ACTIVATE can guarantee a level of quality on the returned options that is not possible to do with CHOOSE.
4.3 [Uncaptioned image]
The CREATE task represents decisions on assembling, synthesizing, or generating new information. This decision represents the creation of knowledge or insights[9, 18, 40, 58], or more tangibly, reports[19], new data transformations[39], etc. In contrast with the other tasks in our typology, CREATE can take any number of options (including none) and return options with new information or modifications, new options (e.g., adding to the set of options or new features to existing options), or returning a set of options of a completely different type (illustrated in the right panel of Figure A Typology of Decision-Making Tasks for Visualization).
Consider again the decision to buy a car. Rather than deciding on an existing set of cars, the CREATE decision allows for three possibilities: (1) the addition of a new car that is not in the current set of options, (2) the modification or inclusion of features such as adding whether each car has advanced driver assist technologies, or (3) synthesizing new information based on the options, like creating a custom car where the initial set of options are components of a car (e.g., body, wheels, software capabilities). The last possibility is perhaps the most impactful, as it describes the case where the evaluation in CREATE returns a new type of option than the initial set of options. To further illustrate (3), consider the more complex decision-making problem of emergency planners deciding who to send evacuation orders to due to a wild fire. An example CREATE decision in this case evaluates predictions of the fire spread (the initial set of options), and returns a set of possible households to evacuate (a new set of options). In this case, the evaluation consists of using one type of information to produce another type.
The CREATE task is often considered a precursor to decisions in the literature. In Simon’s framework, creating alternative courses of action by evaluating their consequences falls under the design stage [61, 62]. However, in agreement with Fischhoff et al. [30], we argue that the CREATE task is more than just a precursor to a decision– it is itself a fundamental decision that can be the ultimate goal of a decision support tool. For instance, generating a report that synthesizes insights while using a tool exemplifies how the CREATE task serves as an end decision goal.
While we can make some guarantees on the quantity or quality of returned options in CHOOSE and ACTIVATE, we cannot do so for CREATE. If the initial options do not meet a set of criteria, this can propagate through the CREATE evaluation and impact the quality of the returned options. Furthermore, even if the initial options do meet a set of criteria, the returned options may require a different set of criteria in order to assess their quality. Constraints on the quantity of returned options from the CREATE evaluation could arise from initial options that do not meet a set of criteria. However, the occurrence of inspiration, innovation, or randomness in the CREATE evaluation can allow an unbounded number of returned options. The quality and quantity assertions for CHOOSE and ACTIVATE are possible because we can reason directly on the initial options in relation to their evaluations. While we can also reason about the impact of the initial options on the CREATE task, the CREATE evaluation is less bounded in that there is no prescribed way to use the initial options. This leads to additional uncertainty in the quality and quantity of returned options than what is present in CHOOSE and ACTIVATE.
4.4 Discussion of Definitions
The three decision tasks are distinguished by how they evaluate information, or how a decision is made. The same decision goal (e.g., buying a car) is achievable with any decision task if the appropriate information (e.g., the cars, their features, and the criteria) is available and the evaluation is possible. For example, a user could prefer an ACTIVATE task because evaluating options individually by their features is easier than assessing options against one another. Alternatively, another user may prefer to execute a CHOOSE task because assessing an option in the context of the available options is helpful. Or perhaps yet another user prefers to execute a CREATE task because they prefer to devise their own option while drawing inspiration from the available options. Although we use clear language distinguishing options from features and criteria, the decision tasks can operate on any of these entities. For example, as part of the decision goal of buying a car, we could CREATE the criteria that are then used to evaluate cars. The decision tasks allow us to describe different approaches in decision-making between different users or algorithms, as well as changes in approach that can occur due to evaluation difficulties or information availability.
The decision tasks in our typology are independent from but synergistic with the intent (task framing), judgement (valuation of information), or analytical actions that can occur while executing the task evaluation. Specifically, these facets of decision-making do not change the evaluation mechanics of the decision tasks, but they can impact the outcome of a decision task. For example, an ACTIVATE task does not depend on whether a user executed it with the intent to include the best options or to exclude the worst options, as either intent can be achieved with an ACTIVATE evaluation. However, the intent may influence how each option is assessed and therefore the outcome of the ACTIVATE task. This distinction is further highlighted when considering the judgement and analysis of information. An anchoring bias in a CHOOSE task influences how later options are evaluated, but it does not change the details of the evaluation. Whether we use sorting or correlation as an analytical technique in CHOOSE only impacts which options are returned. Our decision tasks describe the unique functions of how decisions are made, while these other facets of decision-making help explain potential differences in decision outcomes.
The distinction on how information is processed to make a decision, rather than what influences the outcome of a decision, helps ensure our typology is simple, generalizable, and accessible (design goals G4-7). The flexibility to describe a decision goal using any decision task, or combination of decisions tasks, hints at important properties of our typology, namely the composability and hierarchical structure of the tasks, which we discuss in the next section.
5 Using the Typology for Decision-Making Problems
In isolation, each decision task already provides new ways to express decisions in relevant ways for visualization. However, many decision problems are complex and involve multiple related decisions[61]. The decision tasks in our typology can be combined to represent these larger decision-making problems at various levels of detail because the tasks are composable and hierarchical. Furthermore, we can use a diagram to illustrate these properties and the inherent connections between the tasks within our typology.
5.1 Properties that help describe decision problems
Consider again the example decision of wildfire evacuation: CHOOSE an evacuation location for residents. We could decompose this CHOOSE task into several sub-decision tasks like a CREATE task that decides a possible set of locations given predictions of where the fire may spread, and an ACTIVATE task that decides which locations are far enough from the fire zone yet still in the vicinity for evacuated residents, and enough capacity for everyone. The ability to break down the large CHOOSE task into sub-tasks illustrates how smaller decisions help resolve the overall decision. This example illustrates two properties of our typology:
Composability: The sub-decision tasks are interconnected and describe a flow among decisions. These sub-decisions are said to be composable because the output of one task becomes the input to the next (design goals G1,G3). In our example, CREATE decides potential options and ACTIVATE then decides which options meet given criteria. Note that the execution of CREATE must occur before the ACTIVATE decision because the information output by CREATE is the information evaluated in ACTIVATE. The property of composability is important for describing visualization tasks in general, as demonstrated by the task typology by Brehmer et al.[13]. Composability allows us to describe the relationship between decisions, which in turn helps us understand how information informs decisions (e.g., we know the set of options output by the ACTIVATE decision are a good set to CHOOSE from) and is transformed by decisions (e.g., predictions of wildfire spread into potential evacuation locations).
Hierarchical: The decomposition of the CHOOSE task in the above example into sub-tasks implicitly defines a hierarchical structure of decision tasks (design goal G2). They provide a high-level abstraction of a decision task, while detailing the step-by-step decision-making processes at lower levels. The hierarchies indicate which lower-level decisions are needed to execute higher-level decisions. Furthermore, it enables us to describe situations when multiple disparate decisions make-up a larger more complex decision.
5.2 Diagramming Decision-Making Problems
Figure 1: A diagram of the example wildfire evacuation decision-making problem. The CREATE task has an arrow into the ACTIVATE task, indicating that the CREATE task must occur before the ACTIVATE task. The ACTIVATE task has a self-loop, which indicates the possibility that this task is executed more than once. Lastly, the ACTIVATE task has an arrow into the CHOOSE task, indicating that the locations decided on in the ACTIVATE task are the input to the CHOOSE decision.
We use a directed graph, referred to as a “diagram,” to illustrate decision problems represented in our typology. An example for the wildfire evacuation decision is shown in Figure 1. In this diagram, decision tasks are depicted using distinct shapes: a circle for CHOOSE, a square for ACTIVATE, and a triangle for CREATE. These symbols are supplemented with free-form annotations in domain-specific language to provide details about the decisions being made. Arrows indicate the flow of information between decisions, with annotations on the edges specifying what information is being transferred. Additionally, the “LEVEL X” notation is employed within the case studies to represent the hierarchy of decisions: a LEVEL 1 decision corresponds to the overall goal, a LEVEL 2 decision supports the LEVEL 1 goal, a LEVEL 3 decision supports a LEVEL 2 decision, and so on.
@@ 6 Case Studies
To illustrate further our typology and its properties, we present four case studies where our typology is used to describe user decisions in the context of decision-support tools.
@@ 6.2 Case Study 2: Umbra
Level 1: At the root level, Umbra supports the ACTIVATE decision by evaluating whether the modifications to a sanitized dataset are acceptable, based on its remaining utility and the level of privacy achieved.
@@ 6.3 Case Study 3: Centaurus
Level 2: The Level 1 ACTIVATE decision is supported by additional decisions – an ACTIVATE to select relevant articles based on search terms, a CHOOSE to choose articles for deeper analysis, and another ACTIVATE to search for similar articles after a user has made a decision about a potential article to investigate further.
@@ Discussion:
The WireVis case study diverges from previous examples in that the root decision is not to simply choose or filter out potential options. Instead, it involves the creation of insights assembled from evidence gathered in the tool. Interestingly, the CREATE decision while important in the decision-making problem addressed by this tool, is not directly supported by the visualization system.
@@ 7 Evaluation
To further validate our typology, we recruited nine visualization experts who built decision-support tools in the past to evaluate our typology and identify its strengths and weaknesses. Three of our participants are assistant professors, four are scientists at national laboratories, one is a senior visualization developer at a market research company, and one is a PhD student in visualization. The interviews were semi-structured, guided by a set of pre-determined open-ended questions (please see figure 6) while allowing room for deeper exploration of particular topics as they emerged. Specifically, our questions examined three dimensions proposed by Beaudouin-Lafon[10]: (1) descriptive power: “the ability to describe a significant range of existing [designs]”, (2) evaluative power: “the ability to help assess multiple design alternatives”, and (3) generative power: “the ability to help designers create new designs.” Each interview lasted approximately one hour and was recorded with the participants’ consent.
During each session, we introduced the typology and its properties using the car buying example, as detailed in Sections 4 and 5. We encouraged participants to ask questions and seek clarifications to ensure they fully understood the typology and its properties. Following this, we conducted a collaborative sketching session in shared presentation slides where participants described the decision-making tasks supported by their choice of a visualization system they developed using the typology. As the participants articulated the decision-making tasks and processes supported by their tool, we sketched diagrams (see Section 5.2). Participants could then comment on and refine the diagram until it accurately represented the decision flow of their users. For reference, the slides and diagrams from the interviews are available in the Supplemental Material.
At the conclusion of each interview, we invited participants to provide additional feedback on the criteria used by Brehmer et al.[13] when evaluating their typology: correctness (subsection 7.1), utility in the design process (subsection 7.2), and comparison to other visualization typologies (subsection 7.3). We also asked them to share suggestions for improving the typology both during the interview and afterward via email if they had further insights. We discuss the limitations of the typology gathered from these interviews in more depth in Section 9.
@@ 7.1.1 Completeness
In evaluating the correctness of our typology of decision-making tasks for visualization, we first assessed its completeness to determine if there were any missing decision tasks in our typology that could limit its ability to capture all aspects of decision-making. We adopted Gleicher et al.’s definition of completeness by replacing “designs” with “decision-making problems”, that is, the typology needs to cover all decision-making problems we have encountered [32]. The participants in our study found that our typology was indeed comprehensive, allowing them to represent all decisions within the problem domain. As noted by one of the participants: “I think [the typology] is a pretty good set of high level operations. […] It took me a little bit to really clarify the distinction between activate and choose. But once I got there, it makes a lot of sense. I don’t know if I can really think of another type of operation that I felt was missing. It seemed, at least in the case of my problem, a good set [of operations] to describe what’s going on.” Other participants expressed similar sentiments. Overall, all of the participants were able to describe their decision-making problems using our typology, and none of them identified any missing decision tasks or fundamental operations in our typology.
Interestingly, a few participants incorporated additional annotations when describing their decision-support tools using our typology. These annotations were used to detail the contexts of these participants’ decision-making problems, including temporal information (e.g., time duration for how long a task would take or how frequently a task needs to be performed), roles (e.g., who in an organizational hierarchy should make the decision, or in a collaborative setting, which collaborator is responsible for which task), and decision styles (e.g., rational, conceptual, analytical, etc.). Such contextual annotations complement our typology and suggest an interesting future extension of our work. We explore these concepts further in Sections 9 and 10.
@@ 7.1.2 Expressiveness
During this study, it became clear that our typology is also expressive, meaning it can be easily used to describe decision-making scenarios. All participants expressed their proficiency in composing and decomposing decisions with the typology, enabling them to construct cohesive representations of the problems their respective visualization tools addressed. They noted that the typology captured the decision-making processes inherent in problem domains and facilitated the visualization of user interactions with decision-support tools. As one of the participants said, “[The use of the typology] really kind of helped highlight some of the complexity around the design goals that went into it. And it’s definitely reflective of conversations that [my team and I] had going into the design of the tool and sort of throughout its life span. So, in that sense, I think it’s very representative.